OpenStudy (lukecrayonz):

@jim_thompson5910 help with trig identities real quick?

OpenStudy (lukecrayonz):

cot x sec^4 x=cot x + 2 tan x + tan^3 x

OpenStudy (anonymous):

sec^2=1+tan^2

OpenStudy (lukecrayonz):

Steps?:O

OpenStudy (anonymous):

Do you need help?

OpenStudy (lukecrayonz):

Yes

OpenStudy (anonymous):

That should help.

OpenStudy (lukecrayonz):

Uhh what site is that.. :O

OpenStudy (lukecrayonz):

@zordoloom

jimthompson5910 (jim_thompson5910):

cot x sec^4 x=cot x + 2 tan x + tan^3 x cot x sec^4 x=1/tan x + 2 tan x + tan^3 x cot x sec^4 x=1/tan x + 2 tan^2 x/tan x + tan^3 x cot x sec^4 x=(1 + 2 tan^2 x)/tan x + tan^3 x cot x sec^4 x=(1 + 2 tan^2 x)/tan x + tan^3 x/1 cot x sec^4 x=(1 + 2 tan^2 x)/tan x + tan^4 x/tan x cot x sec^4 x=(1 + 2 tan^2 x + tan^4 x)/tan x cot x sec^4 x=(1 + 2z + z^2)/tan x ... Let z = tan^2 x cot x sec^4 x=((z+1)^2)/tan x cot x sec^4 x=((tan^2 x+1)^2)/tan x cot x sec^4 x=((sec^2)^2)/tan x cot x sec^4 x=(sec^4 x)/tan x cot x sec^4 x=cot x sec^4 x

OpenStudy (lukecrayonz):

;__; you have different answers?

jimthompson5910 (jim_thompson5910):

there are different ways to do this, but they are both valid

jimthompson5910 (jim_thompson5910):

notice how I'm only manipulating the right side and I'm changing it into the left side

OpenStudy (lukecrayonz):

1/tan x + 2 tan^2 x/tan x + tan^3 x (1 + 2 tan^2 x)/tan x + tan^3 x What allowed you to do that?

jimthompson5910 (jim_thompson5910):

1/tan x and 2 tan^2 x/tan x are fractions with the same denominator (tan x)

jimthompson5910 (jim_thompson5910):

I combined them to get (1 + 2 tan^2 x)/tan x

OpenStudy (lukecrayonz):

Thanks a lot @jim_thompson5910 :D

jimthompson5910 (jim_thompson5910):

you're welcome